# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

### v2

@@ 3.4 Distribution of Similarity Scores
Figure 4 compiles the histograms for all nine similarity metrics. Most distributions are unimodal and right-skewed, especially lexical and semantic measures, indicating that the majority of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents. Structural similarity, by contrast, is more dispersed, reflecting Grokipedia’s simpler sectioning and reference patterns.
@@
Figure 4: Empirical distributions of similarity scores across metrics. Each panel corresponds to one of the nine metrics.
Together, these findings reinforce that Grokipedia functions less as an independent encyclopedic corpus than as an AI-generated echo of Wikipedia, maintaining meaning and tone while attenuating the structural and evidentiary scaffolding that underpins human-curated knowledge.
@@ 4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 1,800 matched article pairs drawn from the 2,000 most-edited English Wikipedia titles. Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.82 and stylistic similarity around 0.83, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels (greater syntactic complexity) but lower lexical diversity and reference density. This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
The similarity distributions reveal further nuance. While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are more dispersed, suggesting heterogeneous modes of generation. Some Grokipedia pages closely mirror Wikipedia’s organization and tone; others diverge sharply, producing verbose or restructured narratives. This heterogeneity implies a hybrid authorship logic: part replication, part improvisation.
@@
These results also resonate with earlier linguistic analyses of Wikipedia itself. Yasseri, Kornai, and Kertész [2012a] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts. Our findings parallel this pattern in reverse, Grokipedia’s language inflates sentence length and readability grade without a proportional increase in lexical variety. In both cases, surface-level stylistic change (simplification or elaboration) alters perceived accessibility more than underlying vocabulary structure. Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
@@
Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [2, 7, 24]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21]. Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density. In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution. Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
@@
As argued by Yasseri [24], Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable. Grokipedia inverts this model. Its authorship is singular, automated, and invisible, bias becomes latent, hidden within model weights and unseen editorial heuristics. This creates what might be termed an epistemic opacity paradox: Grokipedia appears neutral because it lacks human editors, yet its underlying generative logic is uninspectable and unaccountable. The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
@@
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 2,000 most-edited pages, of which around 1,800 have counterparts on Grokipedia. This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [26]. Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while Grokipedia’s generative parameters may change with future model retraining. Finally, Grokipedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.

### v3

@@ 3.4 Distribution of Similarity Scores
Figure 4 compiles the histograms for all eight similarity metrics. Most distributions are binomial, indicating that a portion of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents while the others are significantly distinct.
@@
Figure 4: Empirical distributions of similarity scores across metrics. Each panel corresponds to one of the eight metrics. (a) TF–IDF cosine, (b) Jaccard (unigram) (c) Overlap (1-gram) (d) Overlap (2-gram) (e) Semantic cosine (g) BERTScore (F1) (f) Stylistic similarity
To reduce this multidimensional space, we conducted a principal components analysis (PCA). The first component explained the vast majority of variance, with all metrics loading in the expected direction, indicating that the measures capture a common underlying dimension of article similarity. We therefore use this component as a combined similarity score in subsequent analyses. Figure 5 shows the distribution of the combined similarity score.
Figure 5: The distribution of the combined similarity scores.
The distribution in Figure 5 is distinctly bimodal, suggesting the presence of two substantive groups of article pairs: one in which Grokipedia and Wikipedia differ substantially, and another in which the two versions are highly similar.
3.5 Political Bias
Another dimension of article change concerns shifts in political orientation. To quantify this, we examined the references cited in each article and assigned each source a bias score using a publicly available dataset (see Methods for details). For each article pair, we computed the bias shift as the difference between the average bias score of the Grokipedia article and that of its Wikipedia counterpart. Positive values indicate that the Grokipedia version relies more heavily on right-leaning sources, whereas negative values reflect a shift toward left-leaning sources.
Figure 6 plots the bias shift for each article pair against its combined similarity score. As shown in the figure, articles that differ more substantially across platforms are markedly more likely to exhibit a right-leaning shift, whereas highly similar articles tend to cluster near zero, indicating minimal ideological movement.
Figure 6: Bias shift vs. article similarity. Each point represents an article pair. The $x$-axis shows the combined similarity score between the Wikipedia and Grokipedia versions (higher values indicate greater similarity), and the $y$-axis shows the political bias shift (Grokipedia - Wikipedia). Positive values indicate a shift toward right-leaning sources, while negative values indicate a shift toward left-leaning sources. Colored markers denote the centres of mass for the low-similarity and high-similarity groups. As similarity increases, bias shifts converge toward zero; when similarity is low, the magnitude and dispersion of bias shifts increase.
To further examine differences in the sources cited within the least similar article pairs, we focus on the subset of articles with a negative combined similarity score. Figure 5 displays the top 40% most frequently cited news sources within this low-similarity group for each platform. Each wedge is colour-coded according to the ideological bias of the outlet. As shown in the figure, the weighted average bias score of Grokipedia articles in this subset is approximately 25% left-leaning, compared with 31% left-leaning for Wikipedia. This difference indicates a relative rightward shift in the sourcing of the least similar Grokipedia articles, consistent with the pattern observed in the bias–similarity relationship above.
Figure 7: Top news sources among low-similarity articles (negative combined similarity score). Wedges show each outlet’s share of all mentions within this subset; colors reflect the outlet’s political-leaning score from the public News Media Bias and Factuality dataset. [^5] https://huggingface.co/datasets/sergioburdisso/news_media_bias_and_factuality Titles report the overall (news-only) mention-weighted bias for each platform (normalized to $\pm 100\%$).
3.6 Comparison within Topics
Because these articles cover a diverse set of subject areas, we additionally classified each article using GPT-5 (see Methods for details). This allows us to compare similarity and bias shift not only across individual articles but also across broader thematic domains.
Figure 8 summarizes these findings by plotting, for each topical category, the average similarity between Wikipedia and Grokipedia articles together with the corresponding average shift in political bias. As shown in Figure 8(a), the largest cross-platform differences in similarity appear in articles related to politics, geography, history, business, and religion. In parallel, as shown in Figure 8(b), the strongest rightward shifts in source bias occur in articles on religion, history, languages, and business, indicating that ideological divergence is especially pronounced in these domains.
Figure 8: Topic-level comparison of similarity and political bias shift between Wikipedia and Grokipedia articles. (a) Mean combined similarity scores between Wikipedia and Grokipedia articles within each topic, with standard-error bars. (b) Mean political bias shift (Grokipedia – Wikipedia) for articles in each topic, indicating whether Grokipedia content tends to lean more left or right relative to Wikipedia. Error bars show standard errors within each topic.
@@ 4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 17,700 matched article pairs drawn from the 20,000 most-edited English Wikipedia titles. Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.86 and stylistic similarity around 0.88, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels and lexical diversity (greater syntactic complexity), but reference density. This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
An important distinction between the two platforms lies in their use of supporting media and internal link structure. Unlike Wikipedia, the current version of Grokipedia includes no images and does not offer linking to related pages within the platform. These omissions substantially reduce navigational affordances and make Grokipedia articles harder to scan, explore, and contextualize.
The similarity distributions reveal additional nuance. The bimodal pattern across measures indicates that while some Grokipedia pages closely mirror Wikipedia’s organisation and tone, others diverge sharply. This heterogeneity suggests a hybrid authorship logic: part replication, part improvisation. In some cases, Grok appears to reproduce Wikipedia’s structure with high fidelity; in others, it departs substantially from the source material, generating content that is stylistically and semantically distinct.
@@
These results also resonate with earlier linguistic analyses of Wikipedia itself. Yasseri, Kornai, and Kertész [yasseri2012complexity] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts. Our findings parallel this pattern in reverse: Grokipedia’s language inflates sentence length and readability grade. In both cases, surface-level stylistic change (simplification or elaboration) alters perceived accessibility more than underlying vocabulary structure. Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
@@
Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [apnews2025, businessinsider2025, yasseri2025conversation]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [rogers2025wired, oremus2025, hartwelle2025, wong2025]. Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density. In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution. Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
@@
As argued by yasseri2025conversation, Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable. Grokipedia inverts this model. Its authorship is singular, automated, and invisible, bias becomes latent, hidden within model weights and unseen editorial heuristics. This creates what might be termed an epistemic opacity paradox: Grokipedia appears neutral because it lacks human editors, yet its underlying generative logic is uninspectable and unaccountable. The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
Complementing these textual and stylistic findings, our reference-level analysis (Figures 65) reveals that the magnitude and direction of political bias shifts between Grokipedia and Wikipedia articles are strongly related to article similarity. When Grokipedia reproduces Wikipedia closely (high similarity), the resulting entries show little or no bias shift. However, low-similarity articles—those most substantially rewritten—display both greater variance and a systematic rightward tendency in bias shift. This pattern suggests that editorial divergence in Grokipedia is concentrated precisely where generative rewriting occurs.
The corresponding breakdown of reference sources further supports this interpretation. In the low-similarity subset, citation networks are dominated by a narrower set of news outlets with identifiable political leanings, as measured by the News Media Bias and Factuality dataset [sanchez2024mapping]. Whereas Wikipedia’s reference base remains diverse and balanced, Grokipedia’s alternative sourcing emphasizes ideologically marked or partisan media domains. Together, these results indicate that Grokipedia’s generative process may amplify latent source bias when departing from human-authored baselines, substituting breadth and attribution with ideologically skewed curation.
The topic-level analysis reveals that cross-platform divergence is not evenly distributed but concentrated in specific domains. Articles on politics, geography, history, business, and religion show the lowest similarity between Grokipedia and Wikipedia, suggesting that the model’s generative behavior departs most from human-edited content precisely in areas where public narratives are contested or where factual structures are complex. Notably, some of these domains exhibit the strongest rightward shift in cited sources, too. This alignment between low similarity and increased ideological skew suggests that Grokipedia’s deviations are not merely stylistic but often substantively directional.
@@
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 20,000 most-edited pages, of which around 17,700 have counterparts on Grokipedia. This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [yasseri2012dynamics]. Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while Grokipedia’s generative parameters may change with future model retraining. Finally, Grokipedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.